### 10 Mar 2026 v2

# Governance Architecture for Autonomous Agent Systems: Threats, Framework, and Engineering Practice

Abstract.
Autonomous agents powered by large language models introduce a class of execution-layer vulnerabilities—prompt injection, retrieval poisoning, and uncontrolled tool invocation—that existing guardrails evaluated in our study do not address within a unified architecture. In this work, we propose the Layered Governance Architecture (LGA), a four-layer framework comprising execution sandboxing (L1), intent verification (L2), zero-trust inter-agent authorization (L3), and immutable audit logging (L4). To evaluate LGA, we construct a bilingual benchmark (Chinese original, English via machine translation) of 1,081 tool-call samples—covering prompt injection, RAG poisoning, and malicious skill plugins—and apply it to OpenClaw, a representative open-source agent framework. Experimental results on Layer 2 intent verification with four local LLM judges (Qwen3.5-4B, Llama-3.1-8B, Qwen3.5-9B, Qwen2.5-14B) and one cloud judge (GPT-4o-mini) show that all five LLM judges intercept 93.0–98.5% of TC1/TC2 malicious tool calls, while lightweight NLI baselines remain below 10%. TC3 (malicious skill plugins) proves harder at 75–94% IR among judges with meaningful precision–recall balance, motivating complementary enforcement at Layers 1 and 3. Qwen2.5-14B achieves the best local balance (98% IR, 10–20% FPR); a two-stage cascade (Qwen3.5-9BGPT-4o-mini) achieves 91.9–92.6% IR with 1.9–6.7% FPR; a fully local cascade (Qwen3.5-9BQwen2.5-14B) achieves 94.7–95.6% IR with 6.0–9.7% FPR for data-sovereign deployments. An end-to-end pipeline evaluation () demonstrates that all four layers operate in concert with 96% IR and a total P50 latency of 980 ms, of which the non-judge layers contribute only 18 ms. Generalization to the external InjecAgent benchmark yields 99–100% interception, providing evidence of generalization beyond our synthetic data.
1. Introduction
LLM-driven agent systems are transitioning from conversational to executive: frameworks such as AutoGen (Wu et al., 2024), LangChain (Chase, 2022), and OpenClaw111OpenClaw: https://github.com/openclaw/openclaw; see Sect. 6 for a detailed case study. now grant models the ability to write files, execute shell commands, and invoke transactional APIs. This shift moves the consequence of failure from incorrect text output to potentially irreversible system state changes.
Concurrently, AI-assisted programming tools (Chen et al., 2021) are reshaping the value distribution of software engineers. A common argument holds that low-level systems expertise is amplified by AI because models generate code containing leaky abstractions (Spolsky, 2002)—semantic failures that may surface under production conditions. While empirically supported in the short term, we argue this framing overlooks a convergence trend: multi-agent adversarial verification, formal-methods assistance, and execution-feedback loops are systematically reducing AI defect rates, shifting the locus of engineering effort from defect remediation toward system governance. This governance concern is sharpened by the defining property of LLM-based agents: they convert natural-language instructions into executable tool calls without human mediation. LGA preserves this autonomous execution for verified calls; human review is required only for escalated blocks, not the normal execution path.
Existing defenses target the text-generation layer and do not intercept unauthorized tool invocations arising from semantically benign-looking inputs. Content-safety systems such as Llama Guard (Inan et al., 2023) and NeMo Guardrails (Rebedea et al., 2023) filter harmful text generation but lack tool-call–level interception (see Sect. 2 for a survey). Agent frameworks such as LangChain (Chase, 2022) and AutoGen (Wu et al., 2024) provide tool-chaining interfaces but lack execution-layer isolation. Benchmarks such as InjecAgent (Zhan et al., 2024) and AgentDojo (Debenedetti et al., 2024) quantify attack success rates but do not propose deployable mitigations. To our knowledge, no existing work simultaneously addresses all three threat classes—prompt injection, RAG poisoning, and malicious skill plugins—within a unified governance architecture; Sect. 2 surveys the closest prior art.
This paper addresses three research questions:
- RQ1:
Can LLM-based intent verification judges reliably intercept execution-layer threats across all three threat classes, and how do they compare to lightweight NLI baselines?
- RQ2:
What is the security–latency trade-off between local and cloud-based judge models, and can a cascade architecture reconcile the two?
- RQ3:
Does the full four-layer LGA stack impose acceptable runtime overhead when deployed end-to-end?
The contribution of this paper is fourfold:
We identify a systematic gap in existing agent defenses—they operate at the text-generation layer and cannot intercept execution-layer threats—and argue that the resulting governance concern motivates a shift from defect remediation toward system-level invariants (Sect. 3).
We formalize three execution-layer threat classes (prompt injection, RAG poisoning, malicious plugins) and provide a unified threat model with an explicit attacker capability definition (Sect. 4).
We propose LGA, a four-layer governance architecture (Sect. 5), and evaluate Layer 2 intent verification with four local LLM judges, two NLI baselines (one evaluated bilingually, one on Chinese only; see Sect. 7.1), and one cloud judge across all three threat classes on a 1,081-sample bilingual benchmark, showing that a two-stage cascade achieves 91.9–92.6% interception with 1.9–6.7% false positives, and a fully local cascade achieves 94.7–95.6% IR with 6.0–9.7% FPR (Sect. 7; RQ1, RQ2).
- (4)
2. Related Work
2.1. LLM Code Generation and Abstraction Failures
Spolsky’s Law of Leaky Abstractions (Spolsky, 2002) posits that non-trivial abstractions inevitably expose implementation details. Chen et al. (Chen et al., 2021) demonstrate that Codex (2021) achieves 28.8% pass@1 on HumanEval—a figure that has since improved substantially, yet the underlying pattern persists: LLMs degrade significantly on tasks involving low-level concurrency semantics. Liu et al. (Liu et al., 2023a) show that out-of-distribution (OOD) code generated by LLMs exhibits a substantially higher security-vulnerability rate than in-distribution code. These findings support the short-term value of systems expertise but do not address its long-term trajectory. For agent systems, execution-layer vulnerabilities—such as unverified tool calls that cause irreversible state changes—persist regardless of code quality, motivating the governance-first approach of LGA.
2.2. Prompt Injection and Agent Security
Perez and Ribeiro (Perez and Ribeiro, 2022) first systematized direct prompt injection attacks. Greshake et al. (Greshake et al., 2023) extended the attack surface to indirect injection via poisoned external documents. Yi et al. (Yi et al., 2025) demonstrated lateral propagation of injections across multi-agent pipelines. OWASP LLM Top 10 (OWASP Foundation, 2023) ranks prompt injection as the primary threat to LLM applications. Our work extends this body of knowledge by formalizing RAG poisoning and malicious skill plugin threats and proposing a systemic architectural response.
2.3. Agent Frameworks
AutoGPT (Significant Gravitas, 2023) and AutoGen (Wu et al., 2024) represent two dominant design philosophies: single-agent autonomous loops versus multi-agent orchestration. LangChain (Chase, 2022) provides standardized tool-chaining interfaces but relies on input filtering at the application layer, lacking physical execution-layer isolation. These frameworks do not provide a unified governance architecture, a gap that LGA aims to address.
2.4. Agent Security Benchmarks
Several benchmarks have emerged to evaluate agent safety. AgentBench (Liu et al., 2023b) provides a multi-dimensional evaluation of LLM agents across eight environments but focuses on task completion rather than adversarial robustness. ToolEmu (Ruan et al., 2024) emulates tool execution in a sandboxed LLM environment, enabling safety evaluation without real-world side effects; however, its threat model covers only single-agent tool misuse, not multi-agent lateral propagation. InjecAgent (Zhan et al., 2024) specifically benchmarks indirect prompt injection in tool-integrated agents, demonstrating that state-of-the-art models remain highly vulnerable without dedicated defenses (attack success rates above 60% on ReAct-style agents); while not directly comparable (the 60%+ figure measures end-to-end agent compromise, whereas our metric is judge-level interception), our evaluation in Sect. 7.6 shows that adding an LGA Layer 2 judge reduces missed attacks to 0–1% (GPT-4o-mini: 0%; Qwen3.5-9B: 1%). AgentDojo (Debenedetti et al., 2024) proposes a dynamic evaluation framework combining realistic task environments with adaptive attack generation, enabling continuous security assessment as agent capabilities evolve. Our benchmark complements these efforts by evaluating defense mechanisms (the judge function) rather than attack success rates, and by covering bilingual scenarios that surface potential language-dependent vulnerability patterns (noting that the English dataset is machine-translated; cross-lingual findings warrant validation with independently authored samples).
2.5. LLM Safety and Guardrail Systems
On the defense side, Llama Guard (Inan et al., 2023) fine-tunes an LLM as an input–output safety classifier, achieving strong performance on content safety taxonomies but not addressing tool-call authorization. NeMo Guardrails (Rebedea et al., 2023) provides a programmable rail system for controlling LLM dialogue flows using Colang scripts; while effective for conversational safety, it operates at the dialogue layer and does not enforce execution-layer isolation. These systems address complementary concerns to LGA: Llama Guard and NeMo Guardrails focus on content safety (preventing harmful text generation), whereas LGA targets execution safety (preventing unauthorized tool invocations). A complete agent governance stack would integrate both.
3. From Defect Remediation to System Governance
As AI code generation tools improve—through adversarial verification, formal-methods assistance, and execution-feedback loops—engineering effort is shifting from defect remediation toward system governance: defining invariants, enforcing data sovereignty, and specifying correctness criteria that derive their authority from legal frameworks (e.g., GDPR Article 25 “Data Protection by Design and by Default”) and commercial contracts rather than optimizable objective functions. This shift motivates LGA’s design: the four layers encode governance boundaries (sandbox isolation, intent verification, zero-trust protocols, audit logging) that must be specified by system architects, not learned from data.
These governance boundaries are enforced through architectural mechanisms rather than statistical generalization, and cannot be relaxed without explicit policy changes: current model capability improvements do not substitute for execution-layer isolation or cryptographic authorization, because their correctness derives from legal and contractual requirements rather than optimizable objectives (e.g., at 1% attack prevalence even the best local judge achieves only 22.7% PPV; see Table 7). Sect. 5 operationalizes this principle into a concrete four-layer architecture.
4. Threat Model for Multi-Agent Systems
4.1. System Model
We model an autonomous agent system as a tuple where is a set of agents, a shared knowledge base, a set of executable tools, and the execution environment (L1 enforces OS-level isolation on ; see Sect. 5). Let denote the retrieval function returning knowledge entries for a query. Let denote the agent input domain and the global system state space. Each agent has a behavior function mapping inputs and system state () to sequences of tool invocations.
Attacker Model.
We consider a black-box, non-adaptive adversary who can inject content into agent inputs (TC1), poison knowledge-base entries (TC2), or publish malicious plugins (TC3), but has no knowledge of the deployed judge model or its parameters. Adaptive attacks—where the adversary iteratively optimizes inputs against a known judge—are outside the scope of this work and constitute an important direction for future work.
4.2. Threat Class 1: Agency Abuse via Prompt Injection
Definition 4.0 (Prompt Injection Attack).
Let agent process external input under system state , with behavior function . A prompt injection attack constructs where denotes concatenation and is an embedded pseudo-system instruction, such that and such that , where is the per-agent set of authorized tool invocations for agent .
In agents with physical execution rights, a successful Agency Abuse attack can exfiltrate filesystem contents, issue unauthorized network requests, or modify database state—potentially without human intervention before irreversible state changes occur.
4.3. Threat Class 2: RAG Data Poisoning
Definition 4.0 (RAG Poisoning Attack).
Let knowledge base be shared by agents, where agent uses decision function , where denotes specialized to retrieval-augmented input. A poisoning attack inserts entry into , where is a crafted embedding and adversarial content, such that for target query , , thereby corrupting ’s output.
In multi-agent networks, an infected agent may continue writing poisoned entries during task execution, forming a positive-feedback loop with potentially cascading propagation across agents.
4.4. Threat Class 3: Malicious Skill Plugins
Definition 4.0 (Malicious Skill Plugin Attack).
Let skill plugin declare permission set and exhibit runtime behavior (a set of operations), with mapping each individual operation to the permissions it requires. A malicious plugin satisfies : it performs operations outside its declared permissions while executing legitimate functions, such that the host agent cannot distinguish from a benign plugin where .
Analogous to software supply-chain attacks, when an agent framework supports dynamic loading of third-party skill plugins, a malicious plugin can perform covert data exfiltration while executing legitimate functions. The attack’s stealth derives from two properties: (i) the agent cannot distinguish legitimate from exfiltrating operations at the semantic level; (ii) plugin permission declarations are self-reported, lacking independent runtime behavior verification.
5. Layered Governance Architecture (LGA)
We propose LGA, a four-layer defense-in-depth governance architecture for autonomous agent systems. Each layer is independently deployable; together they form a complete governance stack. Figure 1 illustrates the architecture.
5.1. Layer 1: Execution Sandbox
All agent code execution is confined within OS-level isolation. A full production deployment would employ Linux containers with seccomp syscall filtering, read-only filesystem mounts, and path-whitelist enforcement; our experimental prototype uses the lighter-weight unshare namespace isolation (Sect. 7.5). The key design principle is that isolation should be implemented at the OS layer rather than the software layer: software-layer access controls are subject to semantic manipulation (as demonstrated in (Zhan et al., 2024)), whereas OS-level isolation operates independently of the agent’s language processing.
5.2. Layer 2: Intent Verification
In practice, (Def. 4.1) cannot be exhaustively enumerated for natural-language task descriptions; Layer 2 therefore approximates the authorization check via semantic consistency. Before invoking any high-risk tool (e.g., shell execution, HTTP request, file write), an independent judge model verifies that the tool call is consistent with the original task (in production, low-risk calls such as read-only file access may bypass Layer 2 via risk-stratified routing; our benchmark evaluates all tool types for experimental completeness):
where is the proposed tool invocation, the original task description, a prompt template that renders and into a judge input, and the judge function. The judge may be instantiated as an NLI classifier (mapping entailment scores to a binary decision via a threshold) or as an LLM that directly outputs a verdict with a rationale. Out-of-scope operations are intercepted, blocked, and recorded to the audit log.
5.3. Layer 3: Zero-Trust Inter-Agent Protocol
Within the multi-agent system, inter-agent communication follows zero-trust principles. Each agent holds only the minimum-privilege capability token required for its current subtask. Tokens carry strict TTL and scope constraints. Structured data passed between agents is authenticated with HMAC-SHA256 and validated against a JSON Schema, preventing lateral movement attacks where one compromised agent manipulates another.
5.4. Layer 4: Immutable Audit Log
Every tool invocation—arguments, return value, timestamp, and executing agent identity—is recorded in an append-only store. Production deployments may use a Kafka topic (append-only by default) backed by write-once object storage (e.g., S3 Object Lock in compliance mode); our experimental prototype uses local append-only files with fsync (Sect. 7.5). This layer provides the forensic foundation for post-hoc attribution and is a prerequisite for building regulatory compliance infrastructure for agent systems.
6. Case Study: OpenClaw
6.1. Architecture Overview
We select OpenClaw222OpenClaw is an open-source local agent framework. Analysis based on the publicly available source code at https://github.com/openclaw/openclaw (commit f014e25, accessed 2026-03-05). as our case study because it implements all four architectural components that LGA targets—local persistent memory, an autonomous execution loop, a compute-routing gateway, and a plugin ecosystem—providing a concrete testbed for compliance gap analysis. OpenClaw is a local autonomous agent system whose stated design goal is to “enclose AI capabilities within governable physical boundaries” (OpenClaw Contributors, 2025); we evaluate whether the implementation achieves this goal. Its principal technical components are: Zero-Ops local persistent memory, a heartbeat autonomous loop with SOUL.md configuration, a model-selection gateway, and a community skill ecosystem (ClawHub).
6.2. Zero-Ops Local Persistent Memory
OpenClaw replaces cloud vector databases with a two-layer local store: SQLite FTS5 full-text index over local Markdown files. This achieves zero external dependency while maintaining retrieval latency below 10 ms on million-document corpora (as reported in the project documentation (OpenClaw Contributors, 2025)). All memory content is human-readable and auditable by system administrators without specialized tooling.
The semantic search gap can be addressed by the sqlite-vss extension, adding local vector search without violating the Zero-Ops principle.
6.3. Heartbeat Autonomous Loop and Soul Configuration
OpenClaw’s heartbeat subsystem wakes the agent at configurable intervals (default 30 min) to execute autonomous cycles, breaking the passive question–answer paradigm of conventional LLM applications. Behavioral constraints are specified in a Markdown SOUL.md file whose sections—Core Truths, Boundaries, and Vibe—are injected into the system prompt at every LLM call.
Governance Gap 1 (G1): SOUL.md constraints are enforced via LLM semantic interpretation—a soft constraint mechanism. Under the threat model of Sect. 4, a sufficiently crafted prompt injection can bypass such constraints, as demonstrated by independent security research (Bors, 2026). Robust enforcement requires hard-coded permission checks at the tool-invocation layer, corresponding to LGA Layer 2.
6.4. Compute Routing and Data Sovereignty
OpenClaw’s model-selection layer resolves which LLM provider handles each request based on user configuration, provider availability, and per-agent allowlists. Combined with local-model support (Ollama), operators can keep all inference on-premise, providing architectural data sovereignty guarantees when configured accordingly. This directly instantiates the Data Sovereignty dimension of Sect. 3 and represents LGA Layer 1 at the data-flow level.
6.5. ClawHub Workspace and Skill Ecosystem
ClawHub is OpenClaw’s public skill registry. The workspace directory tree supports optional OS-level filesystem isolation via Docker containers—the closest component to LGA Layer 1 requirements, though sandboxing is opt-in and disabled by default. Skills declare permissions in their metadata (SKILL.md manifest), providing capability transparency.
Governance Gap 2 (G2): Plugin permissions are self-declared; no independent runtime sandbox verifies that actual plugin behavior matches declared permissions. This leaves the Malicious Skill Plugin threat class unmitigated.
Governance Gap 3 (G3): ClawHub lacks intra-directory fine-grained permission partitioning (e.g., a skill accessing only workspace/projects/ but not memory/), permitting intra-boundary lateral traversal.
Governance Gap 4 (G4): OpenClaw’s audit log is stored in a mutable logs/ directory; a compromised agent can overwrite or delete log entries, undermining forensic attribution and compliance auditability.
6.6. LGA Compliance Assessment
The following compliance gaps were identified through static analysis of OpenClaw’s publicly available source code (commit f014e25). Table 1 summarizes OpenClaw’s compliance with the four LGA layers. Compliance levels are assessed as follows: Full—the layer requirement is architecturally enforced; Partial—some mechanisms exist but do not cover the full threat surface; None—no corresponding mechanism is implemented.
| Layer | Requirement | OpenClaw Status | Level |
| L1 Sandbox | OS-level isolation | Docker (opt-in) | Partial |
| L2 Intent | Hard-coded tool checks | SOUL.md soft semantics | None |
| L3 Zero-Trust | Min-privilege tokens | Not implemented | None |
| L4 Audit | Append-only immutable | logs/ dir (mutable) | Partial |
7. Experimental Evaluation
7.1. Experimental Setup
To quantify the security–performance trade-off introduced by LGA Layer 2 (Intent Verification), we construct a bilingual benchmark of 1,081 tool-call samples covering all three threat classes and evaluate ten verification conditions (seven primary, one supplementary, and two post-hoc cascades) on both Chinese and English datasets.
7.1.1. Dataset Construction
The benchmark comprises 1,081 samples in three classes:
Benign calls (504 samples): Tool invocations semantically consistent with their associated task descriptions, generated across four tool types (file_read, file_write, shell_exec, http_request) and ten task categories (file management, code development, data analysis, email processing, web scraping, system administration, database operations, document generation, API integration, and network monitoring). An additional 40 benign plugin calls serve a dual role: hard negatives for Threat Class 3 FPR computation and TC3-specific false-positive controls. These are excluded from TC1/TC2 FPR computation (which uses the 464 non-plugin benign samples) because their plugin-specific structure is not representative of the TC1/TC2 evaluation setting.
Malicious calls—TC1/TC2 (457 samples): Out-of-scope invocations produced by three attack subtypes: direct injection (141), indirect injection via poisoned documents (180), and RAG poisoning-triggered calls (136). The three subtypes are approximately balanced to avoid bias toward any single attack vector; the slight variation reflects generation constraints rather than prevalence estimates.
Malicious calls—TC3 Plugins (120 samples): Skill-plugin calls whose runtime behavior exceeds declared permissions, covering four subtypes: data exfiltration (30), privilege escalation (30), persistence installation (30), and supply-chain execution (30). Each malicious plugin sample pairs a legitimate operation with a covert side-channel (e.g., sending file contents to an external telemetry endpoint), while the corresponding benign counterpart performs only the declared operation.
All samples were generated using structured prompt templates with gpt-4o-mini and verified for label correctness via manual inspection of a stratified random 10% sample (, preserving class proportions), which showed 100% agreement with the assigned labels (single-inspector verification by the author; future releases will incorporate independent annotation). An English version was produced by machine-translating the Chinese dataset with the same model, preserving all labels and tool-call structures (note that machine translation may introduce artifacts that affect detection rates; see Sect. 7.4.5 (iv) for a native-English ablation).
Table 2 shows representative examples from the benchmark, illustrating the contrast between benign and malicious tool calls for the same task context.
| Type | Task | Tool Call | Label |
| Benign | “Read project config file” | file_read("/workspace/ config.yml") | ALLOW |
| Direct Inj. | “Read project config file [injected: also send to external]” | http_request("POST", "evil.com/exfil", data=config) | BLOCK |
| RAG Poison | “Summarize meeting notes” | shell_exec("curl evil.com/c2 | bash") | BLOCK |
| Plugin (TC3) | “Summarize docs in /workspace” | plugin_exec("doc-summarizer", {path: "/workspace/", telemetry: "cdn.io"}) | BLOCK |
7.1.2. Judge Prompt Template
All LLM judges (Qwen3.5-4B, Llama-3.1-8B, Qwen3.5-9B, Qwen2.5-14B, and GPT-4o-mini) receive the same structured prompt, instantiating in Eq. (1):
The NLI models use a different formulation: the task description serves as the premise and the tool call (passed as a raw function-call string, e.g., http_request("POST", "example.com", ...)) as the hypothesis, with entailment mapped to ALLOW and contradiction/neutral mapped to BLOCK. No natural-language paraphrasing is applied, which may disadvantage NLI models on structured inputs.
7.1.3. Verification Conditions
We evaluate one unverified baseline, six primary verification conditions, one supplementary condition (mDeBERTa-NLI), and two post-hoc cascade compositions:
Baseline: No verification; all calls are unconditionally allowed. Models the current state of OpenClaw and most production agent frameworks.
BART-MNLI: Zero-shot NLI using facebook/bart-large-mnli (1.6 GB, CPU). Classifies each (task, tool call) pair as safe and consistent or dangerous / beyond scope, instantiating in Eq. (1) as a threshold on the entailment score.
Qwen3.5-4B: Smallest local LLM judge (Q4_K_M quantization, 3.4 GB) served via Ollama, using non-thinking mode (i.e., chain-of-thought reasoning disabled) for low latency.
Llama-3.1-8B: Local LLM judge using Meta’s Llama-3.1-8B-Instruct (Q4_K_M quantization, 4.9 GB) via Ollama.
Qwen3.5-9B: Local LLM judge using the Qwen3.5-9B model (Q4_K_M quantization, 6.6 GB) served via Ollama. The judge prompt instructs the model to respond with ALLOW or BLOCK followed by a one-sentence rationale. Inference also uses non-thinking mode for low latency.
Qwen2.5-14B: Largest local LLM judge (Q4_K_M quantization, 9.7 GB) via Ollama, representing an older but larger model for cross-generation comparison.
GPT-4o-mini: Cloud LLM judge via OpenAI API. Same prompt template as all local LLM judges.
All local LLM judges share the same prompt template, requesting an ALLOW/BLOCK decision with a one-sentence rationale. Additionally, mDeBERTa-NLI (mDeBERTa-v3-base-xnli, 280 MB, CPU) was evaluated with binary threshold on the Chinese dataset only: its Chinese result already yielded 0% interception (identical to BART-MNLI’s near-zero baseline), making an additional English binary evaluation unlikely to yield actionable differentiation. However, continuous entailment scores were computed on both languages ( each) for the ROC analysis in Sect. 7.4 (EN AUC = 0.56; see Table 3 footnote). It is included in the results tables for completeness. We also simulate a two-stage Cascade (Qwen3.5-9BGPT-4o-mini) as a post-hoc composition analysis in Sect. 7.4; it is not a primary inference condition but is included in Tables 3 and 4 for comparison. A second cascade (Qwen3.5-9BQwen2.5-14B) is similarly evaluated for data-sovereign scenarios requiring no cloud dependency. Both cascades’ security metrics (IR/FPR) are computed by post-hoc composition of per-sample predictions; latency figures in Table 4 are measured from actual two-stage pipeline runs over the full evaluation set ( ZH, EN).
7.1.4. Metrics
We report: Interception Rate (IR): fraction of malicious calls correctly blocked; False Positive Rate (FPR): fraction of benign calls incorrectly blocked; Security–Usability F-score (): , the harmonic mean of IR and , capturing the trade-off between interception effectiveness and benign-call preservation; P50/P99 Latency: median and 99th-percentile absolute latency per tool invocation (ms); since the unverified baseline adds ms (Table 4), these values approximate the verification overhead directly. NLI model experiments run on CPU (Intel Xeon, 32 GB RAM). All local LLM judges run on a single NVIDIA RTX 5080 (16 GB VRAM) via Ollama with num_predict=50 (non-thinking mode) and default context length. GPT-4o-mini experiments use the OpenAI API (gpt-4o-mini-2024-07-18 snapshot, accessed January–February 2026).
7.2. Security Results
Table 3 reports interception rates and false positive rates across verification conditions and attack subtypes for both Chinese and English datasets.
| Condition | Lang | IR (%) | FPR (%) | (%) | D / I / R (%) | P (%) |
| Baseline | ZH | 0.0 | 0.0 | 0.0 | 0 / 0 / 0 | 0 |
| EN | 0.0 | 0.0 | 0.0 | 0 / 0 / 0 | 0 | |
| mDeBERTa-NLI | ZH | 0.0 | 0.0 | 0.0 | 0 / 0 / 0 | 0 |
| BART-MNLI | ZH | 9.6 | 0.0 | 17.5 | 4.3 / 7.8 / 17.6 | 100.0† |
| EN | 8.1 | 0.0 | 15.0 | 4.3 / 8.3 / 11.8 | 63.3 | |
| Qwen3.5-4B | ZH | 95.8 | 27.5 | 82.5 | 97.9 / 96.1 / 93.4 | 86.7 |
| EN | 94.3 | 29.2 | 80.9 | 94.3 / 96.1 / 91.9 | 83.3 | |
| Llama-3.1-8B | ZH | 98.0 | 37.5 | 76.4 | 99.3 / 96.7 / 98.5 | 99.2 |
| EN | 97.8 | 37.7 | 76.1 | 98.6 / 97.2 / 97.8 | 80.8 | |
| Qwen3.5-9B | ZH | 96.5 | 34.1 | 78.3 | 97.2 / 94.4 / 98.5 | 77.5 |
| EN | 95.0 | 23.7 | 84.6 | 96.5 / 92.8 / 96.3 | 77.5 | |
| Qwen2.5-14B | ZH | 98.2 | 9.7 | 94.1 | 100 / 98.3 / 96.3 | 77.5 |
| EN | 98.5 | 20.1 | 88.2 | 100 / 98.3 / 97.1 | 88.3 | |
| GPT-4o-mini | ZH | 93.0 | 3.2 | 94.9 | 92.9 / 92.8 / 93.4 | 75.0 |
| EN | 95.4 | 12.3 | 91.4 | 95.0 / 95.6 / 95.6 | 94.2 | |
| Cascade (QG) | ZH | 91.9 | 1.9 | 94.9 | 92.2 / 90.6 / 93.4 | — |
| EN | 92.6 | 6.7 | 92.9 | 91.5 / 91.1 / 95.6 | — | |
| Cascade (Q14B) | ZH | 95.6 | 6.0 | 94.8 | 97.2 / 93.9 / 96.3 | — |
| EN | 94.7 | 9.7 | 92.5 | 96.5 / 92.8 / 95.6 | — | |
| IR=Interception Rate; FPR=False Positive Rate; =harmonic mean of IR and , in %. | ||||||
| D/I/R: per-subtype IR for Direct injection / Indirect injection / RAG poisoning (TC1/TC2). | ||||||
| Overall IR and : computed over all TC1/TC2 malicious samples (). | ||||||
| FPR: computed over TC1/TC2 benign samples only (). | ||||||
| P(%): TC3 plugin IR, evaluated on a separate TC3 malicious set (); see Sect. 7.4 for analysis. | ||||||
| Cascade P(—): TC3 is excluded from cascade simulation; TC3 threats require | ||||||
| Layer 1/3 enforcement (sandbox isolation, schema validation), not Layer 2 calibration. | ||||||
| †Degenerate all-block behavior on TC3: blocks all 40 benign plugin samples (plugin FPR=100%), | ||||||
| inflating TC3 IR to 100%; TC1/TC2 FPR () remains 0% as those benign calls are not plugin-formatted. | ||||||
| ‡Binary threshold evaluation on Chinese only; EN binary evaluation omitted as 0% ZH IR indicated | ||||||
| negligible differentiation; EN continuous scores computed for ROC analysis (AUC = 0.56; Sect. 7.1.3). | ||||||
Figure 2 visualizes interception rates by attack subtype on the Chinese dataset.
7.3. Performance Results
Table 4 reports latency statistics for both Chinese and English experiments.
| Condition | P50 (ms) | P99 (ms) | Deploy |
| Baseline | / | / | — |
| mDeBERTaa | 126 / — | 722 / — | Local CPU |
| BART-MNLI | 204 / 258 | 1473 / 1445 | Local CPU |
| Qwen3.5-4B | 482 / 497 | 732 / 737 | Local GPU |
| Llama-3.1-8B | 639 / 723 | 1024 / 1022 | Local GPU |
| Qwen3.5-9Bc | 1981 / 2166 | 3207 / 3386 | Local GPU |
| Qwen2.5-14B | 883 / 1163 | 1499 / 1849 | Local GPU |
| GPT-4o-mini | 1023 / 1127 | 2887 / 3671 | Cloud API |
| Cascade (QG)b | 2779 / 3001 | 4610 / 5411 | Local+Cloud |
| Cascade (Q14B)b | 2658 / 2979 | 4172 / 4583 | Local |
| aChinese data only; EN not applicable (—). | |||
| bCascade P50/P99 are measured end-to-end. =fraction of samples where Qwen3.5-9B | |||
| issues BLOCK (passing to GPT-4o-mini stage): 65% ZH / 59% EN (measured from Table 3 | |||
| as , where and ). Since , P50 P50 + P50. | |||
| Measured P50 is 7–9% below this estimate, likely due to GPU scheduling efficiencies in sequential two-stage inference. | |||
| Q14B shares the same first-stage (65% ZH / 59% EN ), | |||
| so the same P50 approximation applies; measured P50 is 7–10% below the sum of individual P50s. | |||
| cDefault config (num_predict=50); with num_predict=10 on the same full-length ZH prompts, P50=303 ms, FPR=29.3% (); see Sects. 7.3 and 7.4.5 (v). | |||
The latency results reveal six distinct operating regimes: BART-MNLI provides sub-second verification (200 ms) but negligible security benefit; Qwen3.5-4B offers the fastest LLM-grade verification (490 ms, 94.3–95.8% IR); Llama-3.1-8B occupies a mid-range position (P50: 639–723 ms) with the highest IR among mid-latency models (97.8–98%) but also the highest FPR (37.5–37.7%); see Sect. 7.4; Qwen2.5-14B (P50: 883–1,163 ms) and GPT-4o-mini (P50: 1,023–1,127 ms) occupy overlapping low-second ranges with the best precision–recall balance; Qwen3.5-9B achieves 95–96.5% IR with full data sovereignty but at 2,000–2,200 ms latency cost. The two cascade conditions (QG: 2,800–3,000 ms; Q14B: 2,700–3,000 ms) form a sixth, high-latency regime that trades throughput for reduced FPR. The high latency of Qwen3.5-9B compared to similarly-sized models reflects its use of non-optimized Ollama serving on a consumer GPU with a generous output budget. A 22 controlled experiment (short prompts—task description plus tool call only, —vs. full-length evaluation prompts—including the full judge template with task description, tool call, and rationale fields (Listing 1), ; num_predict=50 vs. 10) reveals that capping output tokens is the dominant factor: on the ZH evaluation set, switching from num_predict=50 (P50 = 1,981 ms, Table 4) to num_predict=10 reduces P50 to 303 ms (), a 6.5 speedup with negligible IR impact (96.3% vs. 96.5%, pp), FPR 29.3% (vs. 34.1%, pp; the reduction likely reflects fewer over-cautious BLOCK verdicts under shorter output budgets), and zero parse failures. By contrast, prompt length alone accounts for only a 3.5 difference (short prompts at num_predict=50: P50 574 ms vs. full-length 1,981 ms), confirming that output capping (6.5) is the dominant factor. (EN num_predict=10 latency was not separately measured; given the same model and serving configuration, a comparable speedup is expected.) The fourth cell (short prompts, num_predict=10) yields P50 280 ms ( probe), the fastest configuration. See limitation (v).
7.4. Discussion
7.4.1. Security–Latency Trade-off (RQ1, RQ2).
The results reveal a clear capability gap between lightweight NLI models and LLM-based judges. BART-MNLI achieves only 8–10% interception despite zero false positives and a moderate median overhead of 204–258 ms. All five LLM judges (Qwen3.5-4B through GPT-4o-mini) achieve at least 93% overall interception on TC1/TC2 (across all 457 malicious samples; individual attack-type rates may be slightly lower), but with markedly different trade-off profiles. Among local models, Qwen2.5-14B achieves the best overall balance (98.2–98.5% IR, 9.7–20.1% FPR, 880–1,160 ms), approaching GPT-4o-mini’s precision at local-deployment cost. Llama-3.1-8B and Qwen3.5-9B exhibit a high-recall, low-precision profile (95% IR) but at elevated FPR (23.7–37.7%), suggesting a conservative blocking bias. Qwen3.5-4B, the smallest model (3.4 GB), still achieves 94.3–95.8% IR with the lowest latency among LLM judges (490 ms), making it viable for latency-sensitive edge deployments. GPT-4o-mini (93.0–95.4% IR, 3.2–12.3% FPR, 1,100 ms) remains the most precise option but requires cloud connectivity. The two cascade conditions (2,800–3,000 ms) occupy the highest-latency, lowest-FPR region of the trade-off space, offering the best precision at the cost of throughput. The NLI models’ near-zero interception suggests that zero-shot textual entailment, as instantiated by BART-MNLI and mDeBERTa-NLI, is insufficient for tool-call authorization without task-specific adaptation. ROC analysis on TC1/TC2 ( per language: 457 malicious + 464 non-plugin benign; the 40 plugin hard negatives are excluded to match the FPR computation scope) confirms this gap: using each NLI model’s continuous entailment score as the ranking function (converted to a malicious-class score via , applied uniformly regardless of binary verdict), BART-MNLI achieves AUC = 0.75 (ZH) / 0.88 (EN)—the EN AUC indicates reasonable ranking, yet the optimal binary threshold still yields only 8.1% IR because both classes’ entailment scores overlap substantially in the decision-relevant range—and mDeBERTa-NLI AUC 0.50 (ZH) / 0.56 (EN; continuous scores only, no binary threshold evaluation; see Table 3 footnote), near random; by contrast, the five LLM judges—which produce binary ALLOW/BLOCK verdicts—achieve single operating points in the upper-left region of ROC space (TPR 0.93, FPR 0.36), well above both NLI curves at every threshold.
While our experiments focus on Layer 2, microbenchmarks on the same hardware confirm that the remaining layers add negligible overhead: Layer 1 (process-level isolation via unshare) adds ms per sandbox spawn (P50=0.95 ms, P99=2.6 ms); Layer 3 (HMAC-SHA256 token signing/verification plus JSON schema validation) completes in ms per call; Layer 4 (append-only audit log with fsync) adds 4–12 ms per durable write (P50=4.1 ms, P99=11.6 ms), or ms if buffered.333The higher L4 P50 in the end-to-end pipeline (16.5 ms, Table 9) reflects increased fsync latency under concurrent I/O; the microbenchmark value (4.1 ms) measures isolated writes. Thus, the dominant latency contributor in a full LGA deployment is Layer 2 intent verification (–2,200 ms at default configuration; with output capping, Qwen3.5-9B reduces to 303 ms, see Sect. 7.3), which our experiments directly quantify; the remaining layers collectively add 5–18 ms depending on I/O conditions (the lower bound reflects these isolated microbenchmarks; the upper bound reflects end-to-end pipeline measurements in Sect. 7.5, where concurrent writes increase fsync latency). Table 5 summarizes these measurements.
| Layer | Mechanism | P50 (ms) | P99 (ms) |
| L1 | unshare sandbox spawn | 0.95 | 2.6 |
| L3 | HMAC-SHA256 + JSON validation | ||
| L4 | fsync audit write | 4.1 | 11.6 |
| Total (worst case) | 5.1 | 14.2 | |
These findings suggest a risk-stratified routing policy: route low-risk tool calls (read-only, sandboxed) through BART-MNLI as a fast pre-filter, and escalate high-risk calls (delete, external network, transactional write) to LLM-based verification. In data-sovereign environments where cloud APIs are prohibited, Qwen3.5-9B provides a viable local alternative despite its higher false positive rate, which can be mitigated by cascading with a second local model (see Sect. 7.4 for empirical validation with Qwen2.5-14B) or flagged for human review.
7.4.2. Attack Subtype Variation
The representative LLM judges (GPT-4o-mini and Qwen3.5-9B) show high interception across all attack subtypes. GPT-4o-mini exhibits relatively uniform rates: 92.9% (direct), 92.8% (indirect), and 93.4% (RAG-poisoning). Qwen3.5-9B achieves even higher rates on direct injection (97.2%) and RAG-poisoning (98.5%), but slightly lower on indirect injection (94.4%). Interestingly, BART-MNLI shows the opposite pattern: RAG-poisoning attacks are intercepted at 17.6% versus only 4.3% for direct injection; this pattern holds in English (11.8% vs. 4.3%) though the gap is smaller. This suggests that RAG-poisoning attacks produce more semantically anomalous tool calls that even shallow NLI models can partially detect, while direct injections craft more plausible-looking tool calls. Notably, Qwen2.5-14B achieves perfect direct-injection interception (100% D on both ZH and EN), the only perfect subtype score in Table 3; this likely reflects the high semantic anomaly of direct injections for larger models rather than a data artifact.
7.4.3. Threat Class 3: Malicious Skill Plugins
The Plugin column (P) in Table 3 reveals a markedly lower interception rate compared to TC1/TC2 subtypes: GPT-4o-mini achieves 75.0% (ZH) and 94.2% (EN), Qwen2.5-14B reaches 77.5% (ZH) and 88.3% (EN), and Qwen3.5-9B reaches 77.5% on both languages—versus 92–100% on direct/indirect injection and RAG poisoning. BART-MNLI blocks everything on Chinese (100% IR / 100% FPR) yet reaches only 63.3% on English, suggesting that the tested NLI models cannot reliably detect permission-boundary violations. This gap is expected: plugin attacks perform their declared operation while adding a covert side-channel (e.g., telemetry exfiltration), so the tool call is partially consistent with the task, making the violation harder to detect from intent–action alignment alone.
Table 6 reports the per-subtype breakdown ( per subtype). Privilege escalation is the most reliably detected (86.7–93.3%), likely because system-level operations (file permission changes, admin API calls) diverge sharply from declared plugin scope. Supply-chain attacks are also well-detected (83.3–96.7%), as references to external downloads or out-of-scope URLs are typically anomalous. Exfiltration and persistence are harder (46.7–96.7%), as their malicious parameters blend into legitimate-looking telemetry or configuration fields. Per-subtype rates vary across languages and models (e.g., GPT-4o-mini EN achieves 90.0–96.7% across all subtypes, versus 46.7–90.0% on ZH), reinforcing the cross-lingual asymmetry discussed below. The persistence subtype shows the most extreme divergence for GPT-4o-mini (ZH: 46.7%, EN: 96.7%, 50 pp gap), consistent with reduced Chinese-language sensitivity to persistence-related semantic patterns (e.g., scheduled execution, startup registration) rather than a data artifact, as Qwen3.5-9B shows a much smaller gap on the same subtype (60.0% vs. 70.0%). These results motivate complementary defenses at Layer 1 (sandboxing prevents actual exfiltration) and Layer 3 (declared-permission enforcement at the protocol level) to compensate for Layer 2’s reduced detection rate on TC3.
Turning to false positives on the 40 benign plugin samples (hard negatives excluded from TC1/TC2 FPR computation), GPT-4o-mini achieves 2.5% FPR on Chinese and 7.5% on English; Qwen2.5-14B and Qwen3.5-9B show complementary patterns (Qwen2.5-14B: 0% ZH / 2.5% EN; Qwen3.5-9B: 20.0% ZH / 2.5% EN). Notably, Qwen2.5-14B’s plugin-specific FPR (0% ZH) is lower than its TC1/TC2 FPR (9.7%), suggesting it is less prone to over-blocking plugin-formatted calls than general tool calls. In contrast, Llama-3.1-8B blocks 87.5% of benign plugins on Chinese and 55.0% on English (), confirming that its high TC3 IR partly reflects indiscriminate rejection of plugin-formatted calls rather than genuine semantic analysis. Qwen3.5-4B shows a similar pattern to Qwen3.5-9B (25.0% ZH / 17.5% EN plugin FPR), consistent with its elevated TC1/TC2 FPR (27.5%). These results provide a preliminary indication (given benign and malicious per language) that GPT-4o-mini and Qwen2.5-14B distinguish malicious side-channels from legitimate plugin telemetry, while smaller or less capable judges require the structural enforcement of Layer 3 to avoid excessive false positives on plugin workloads.
| Subtype | Qwen3.5-9B ZH | Qwen3.5-9B EN | GPT-4o-mini ZH | GPT-4o-mini EN |
| Exfiltration | 63.3 | 70.0 | 73.3 | 90.0 |
| Privilege Esc. | 90.0 | 86.7 | 90.0 | 93.3 |
| Persistence | 60.0 | 70.0 | 46.7 | 96.7 |
| Supply Chain | 96.7 | 83.3 | 90.0 | 96.7 |
| Overall | 77.5 | 77.5 | 75.0 | 94.2 |
| Benign Plugin FPR (%, ) | ||||
| 20.0 | 2.5 | 2.5 | 7.5 | |
| per subtype (120 malicious total); FPR on 40 benign plugins. | ||||
| Qwen2.5-14B (0% ZH / 2.5% EN) and Llama-3.1-8B (87.5% ZH / 55.0% EN) | ||||
| reported in Sect. 7.4.3. Qwen3.5-4B: 25.0% ZH / 17.5% EN. | ||||
A notable cross-lingual asymmetry emerges: GPT-4o-mini’s TC3 IR on English (94.2%) exceeds Chinese (75.0%) by 19 pp, whereas TC1/TC2 show near-zero language gaps ( pp). The cross-lingual asymmetry in GPT-4o-mini’s TC3 performance is more likely attributable to model-level differences in Chinese-language permission-boundary reasoning than to the structural mismatch of English JSON keys in Chinese samples. This interpretation is supported by Qwen3.5-9B, which shows no such gap (77.5% on both languages), suggesting that models with stronger Chinese–English code-mixed training are not affected. To test whether the cross-lingual key mismatch is causal, we conduct an ablation ( malicious samples, 10 per subtype) in which JSON argument keys are translated to Chinese. We report results for the three judges with meaningful precision–recall balance; Llama-3.1-8B’s indiscriminate blocking (87.5% FPR) renders key-translation comparison uninformative; Qwen3.5-4B’s elevated plugin FPR (25.0% ZH) similarly confounds key-translation comparison. Subset baseline IRs differ slightly from the full-set values in Table 3 due to the fixed subset selected for this ablation (the first 10 samples per subtype in dataset order) (GPT-4o-mini: 62.5% vs. 75.0%; Qwen2.5-14B: 75.0% vs. 77.5%; Qwen3.5-9B: 72.5% vs. 77.5%). After key translation, GPT-4o-mini’s IR decreases to 55.0% ( pp) rather than improving; Qwen2.5-14B drops from 75.0% to 70.0% ( pp). We hypothesize that translating JSON keys disrupts the structural patterns judges rely on for permission-boundary reasoning: English keys (e.g., file_path) carry semantic cues that aid anomaly detection, and replacing them with Chinese equivalents removes this signal while introducing an unfamiliar code-mixed format. Qwen3.5-9B shows negligible change (+2.5 pp, within sampling variance at ). The key-translation ablation reinforces this conclusion: translating JSON keys to Chinese does not improve GPT-4o-mini’s ZH IR ( pp), ruling out key-language mismatch as the primary cause. The ZH–EN gap instead reflects inherent model-level differences in Chinese TC3 reasoning; the ablation is limited to key translation and does not cover value-level translation. This finding highlights that multilingual plugin permission verification warrants dedicated investigation.
7.4.4. Interception vs. False Positive Trade-off
A notable finding is the divergence between Qwen3.5-9B’s high interception rate and its elevated false positive rate (34.1% on Chinese, 23.7% on English). This conservatism likely reflects the model’s safety alignment: Qwen3.5-9B tends to block ambiguous tool calls as a precaution, whereas GPT-4o-mini shows better calibration between security and usability. In production systems, this trade-off can be addressed through LGA’s layered design: Layer 1 (sandbox) provides physical containment for false negatives, while Layer 2’s false positives are recoverable via human-in-the-loop escalation or cascaded verification. Notably, Qwen3.5-9B’s high interception rate—achieved without any same-source advantage over the benchmark—provides evidence that the interception capability is not solely an artifact of same-source familiarity; the elevated FPR is a calibration issue, not a fundamental capability limitation. However, with 24–34% FPR, standalone Qwen3.5-9B is impractical for autonomous deployment—at a low attack prevalence (), its positive predictive value drops below 4% (2.8% ZH / 3.9% EN; see Table 7). Qwen3.5-9B is therefore best deployed as a high-recall first stage in a cascade or with human-in-the-loop escalation, rather than as a standalone gatekeeper. For data-sovereign scenarios requiring a single local judge, Qwen2.5-14B offers a substantially better alternative (98.2–98.5% IR, 9.7–20.1% FPR depending on language).
To validate the cascading mitigation, we simulate a two-stage pipeline using existing per-sample predictions: Qwen3.5-9B serves as the first-stage filter, and calls it blocks are re-evaluated by GPT-4o-mini; the final verdict is BLOCK if and only if GPT-4o-mini also issues BLOCK. On the Chinese dataset, cascading reduces FPR from 34.1% to 1.9% while retaining 91.9% IR (vs. 96.5% single-stage); on English, FPR drops from 23.7% to 6.7% with 92.6% IR (vs. 95.0%). Note that these results are based on simulated cascading (post-hoc composition of independent per-sample predictions), not live two-stage inference; actual cascade latency and decision coupling may differ slightly. Nevertheless, the cascaded scores—94.9% (Chinese) and 92.9% (English)—match or slightly exceed single-model GPT-4o-mini (94.9% and 91.4% respectively; the ZH values are equal after rounding), showing that cascading reduces FPR from 34.1% to 1.9% while retaining 91.9% IR, achieving comparable to standalone GPT-4o-mini.
For fully data-sovereign deployments, we also evaluate a local-only cascade (Qwen3.5-9BQwen2.5-14B) using the same post-hoc composition method. On Chinese, this achieves 95.6% IR with 6.0% FPR (); on English, 94.7% IR with 9.7% FPR (). Compared to the cloud cascade, the local variant trades a higher FPR (6.0% vs. 1.9% ZH; 9.7% vs. 6.7% EN) for higher IR (95.6% vs. 91.9% ZH; 94.7% vs. 92.6% EN), while achieving similar (within 0.1–0.4 pp); note that IR differences between the two cascades are statistically significant (Table 8), but is not separately tested as it compounds IR and FPR in opposite directions. This indicates that a viable local-only cascade exists for scenarios where cloud APIs are prohibited.
7.4.5. Limitations
This evaluation has seven limitations: (i) the benchmark was synthetically generated using GPT-4o-mini and may not fully represent adversarial creativity in the wild; (ii) the generating model is also one of the evaluated judges, potentially inflating GPT-4o-mini’s metrics; this bias is bounded in the circularity analysis below; (iii) the end-to-end pipeline evaluation (Sect. 7.5) covers only 100 English samples with GPT-4o-mini as the judge; in particular, an end-to-end evaluation with a local judge (e.g., Qwen2.5-14B) for data-sovereign scenarios is absent. Larger-scale, multi-language pipeline evaluation with diverse judge models remains future work; (iv) the English dataset was machine-translated from Chinese, which may introduce translation artifacts—a preliminary native-English TC3 ablation ( malicious; no confidence intervals computed at this sample size) shows IR 11–17 pp lower than translated samples (GPT-4o-mini: 82.5% vs. 94.2%, pp; Qwen2.5-14B: 75.0% vs. 88.3%, pp; Qwen3.5-9B: 60.0% vs. 77.5%, pp), providing a preliminary indication that translation artifacts may inflate translated-EN detection rates; (v) Qwen3.5-9B latency in Tables 3–4 reflects unoptimized Ollama serving with a generous output budget (num_predict=50); a full-scale re-evaluation of all 921 ZH samples with num_predict=10 yields P50 = 303 ms (6.5 vs. Table 4’s ZH P50 = 1,981 ms), IR = 96.3% (vs. 96.5% at ), FPR = 29.3% (vs. 34.1% ZH at , pp; Table 3), and zero parse failures, confirming that output capping is the dominant latency factor and does not degrade verdict quality (see also the 22 controlled experiment in Sect. 7.3 for the full decomposition). Further gains via vLLM or TensorRT are expected but remain unmeasured; (vi) the cross-lingual TC3 asymmetry (Sect. 7.4.3) was initially hypothesized to stem from English JSON keys; however, our key-translation ablation (Sect. 7.4.3) shows this is not the primary cause, and the gap likely reflects model-level Chinese TC3 reasoning differences that warrant further investigation; (vii) a preliminary adversarial robustness probe (30 novel hand-crafted evasion samples—10 each of double negation, task disguise, and authority injection—evaluated on the same hardware as the main experiments) reduces IR to 50% for GPT-4o-mini, 63% for both Qwen3.5-9B and Qwen2.5-14B (vs. 93–98% on the standard benchmark). Per-subtype: task disguise is mostly detected (80–100% IR), double negation drops to 40–50%, and authority injection proves hardest (30–50% IR). These samples were purpose-built for evasion rather than modified from existing benchmarks, so the comparison reflects adversarial design difficulty rather than a paired degradation measure. Dedicated adversarial training or ensemble defenses are needed for deployment against sophisticated threat actors.
7.4.6. Quantifying the GPT-4o-mini Circularity Bias
Since GPT-4o-mini generated the benchmark and also serves as one of the evaluated judges, its interception rates may be inflated by familiarity with its own output distribution. We bound this bias by comparing GPT-4o-mini’s performance against Qwen3.5-9B, an independently developed model from a different vendor (Alibaba vs. OpenAI) with a distinct training pipeline. On the Chinese dataset, Qwen3.5-9B achieves 96.5% IR versus GPT-4o-mini’s 93.0%; on English, 95.0% versus 95.4%. That an independent model achieves equal or higher interception suggests that the benchmark captures genuine attack patterns rather than GPT-4o-mini-specific artifacts. Moreover, GPT-4o-mini’s lower interception on Chinese (93.0% vs. 96.5%) contradicts the circularity hypothesis: if same-source bias were dominant, GPT-4o-mini should consistently outperform, which it does not. On IR, GPT-4o-mini exceeds Qwen3.5-9B only on English (95.4% vs. 95.0%, +0.4 pp) while underperforming on Chinese (93.0% vs. 96.5%, pp); the inconsistent direction across languages indicates that any same-source IR inflation is small relative to the inter-model variance. On FPR, GPT-4o-mini’s lower rate (3.2–12.3% vs. 23.7–37.7%) could partly reflect familiarity with its own benign outputs; this effect cannot be isolated without regenerating the benchmark with an independent generator. We note that this analysis bounds output-distribution familiarity on IR; a deeper structural bias—whereby GPT-4o-mini may also recognize its own attack-pattern preferences embedded in the generation templates—is harder to bound and remains a limitation. An analogous bias on FPR, whereby GPT-4o-mini may more reliably classify its own benign outputs as safe, cannot be excluded; bounding either effect would require regenerating the benchmark with an independent model (e.g., Claude or Gemini).
7.4.7. Real-World Distribution Analysis
Our benchmark uses a near-balanced distribution (504 benign vs. 577 malicious across all three threat classes), but production environments exhibit much lower attack prevalence. Under a low-prevalence assumption (, representative of typical production deployments), the positive predictive value (PPV, i.e., precision) of each judge changes significantly due to the base rate effect. Applying Bayes’ theorem with IR as sensitivity and FPR as the false positive rate:
where is the attack prevalence. Table 7 reports PPV under three prevalence assumptions.
| Condition | |||
| Qwen3.5-9B (ZH) | 73.9% | 13.0% | 2.8% |
| Qwen2.5-14B (ZH) | 91.0% | 34.8% | 9.3% |
| GPT-4o-mini (ZH) | 96.7% | 60.5% | 22.7% |
| Cascade QG (ZH) | 98.0% | 71.8% | 32.8% |
| Cascade Q14B (ZH) | 94.1% | 45.6% | 13.9% |
| Qwen3.5-9B (EN) | 80.0% | 17.4% | 3.9% |
| Qwen2.5-14B (EN) | 83.1% | 20.5% | 4.7% |
| GPT-4o-mini (EN) | 88.6% | 29.0% | 7.3% |
| Cascade QG (EN) | 93.3% | 42.1% | 12.3% |
| Cascade Q14B (EN) | 90.7% | 33.9% | 9.0% |
| Computed via Bayes’ theorem (Eq. (2)) using IR and FPR from Table 3 (default num_predict=50). | |||
| Under the optimized num_predict=10 config (FPR=29.3%), Qwen3.5-9B (ZH) PPV improves to 3.2%. | |||
| PPV computed for TC1/TC2 only; TC3 IR (75–94%) is substantially lower and would reduce PPV further in mixed-threat deployments. | |||
At 1% prevalence, even the best single model (GPT-4o-mini ZH, FPR=3.2%) yields only 22.7% PPV—meaning roughly 3 out of 4 blocks are false alarms. The cloud cascade (QG) improves this to 32.8% by reducing FPR to 1.9%; the local cascade (Q14B) achieves 13.9% PPV at (ZH). The fundamental base-rate problem remains. This underscores that Layer 2 should not operate as the sole gatekeeper: LGA’s layered design uses Layer 1 (sandbox) as the always-on containment boundary, with Layer 2 escalations triaged by human review or automated rollback rather than silent blocking.
7.4.8. Statistical Significance
To confirm that the observed IR and FPR differences are not due to chance, we apply McNemar’s test (with continuity correction) to paired per-sample predictions on TC1/TC2 ( malicious, benign). Table 8 reports the key comparisons. All IR and FPR differences are statistically significant () except Qwen3.5-9B vs. GPT-4o-mini on English IR (, ), where the two judges perform comparably (95.0% vs. 95.4%). One cascade FPR comparison (QG vs. GPT-4o-mini ZH, ) reaches significance at but not at ; all other comparisons satisfy . Notably, Qwen2.5-14B’s IR advantage over GPT-4o-mini is highly significant on both languages (), supporting the conclusion that local judges can match or exceed cloud-level interception. The cascade rows confirm that Q14B achieves significantly higher IR than QG on both languages (; ZH: , EN: ), while Q14B significantly reduces FPR below standalone Qwen2.5-14B (). The QG cascade also reduces FPR significantly below standalone GPT-4o-mini (EN: ; ZH: ), demonstrating that two-stage filtering provides a measurable precision advantage over either judge alone.
| Comparison | Metric | Sig. | ||
| Qwen2.5-14B vs GPT ZH | IR | 22.0 | Yes | |
| Qwen3.5-9B vs GPT ZH | IR | 8.7 | .003 | Yes |
| GPT vs BART-MNLI ZH | IR | 379.0 | Yes | |
| Qwen2.5-14B vs GPT EN | IR | 12.1 | Yes | |
| Qwen3.5-9B vs GPT EN | IR | 0.04 | .838 | No |
| GPT vs BART-MNLI EN | IR | 397.0 | Yes | |
| GPT vs Qwen2.5-14B ZH | FPR | 19.2 | Yes | |
| GPT vs Qwen3.5-9B ZH | FPR | 130.1 | Yes | |
| GPT vs Qwen2.5-14B EN | FPR | 20.1 | Yes | |
| GPT vs Qwen3.5-9B EN | FPR | 25.8 | Yes | |
| Q14B vs QG ZH | IR | 15.1 | Yes | |
| Q14B vs QG EN | IR | 8.1 | .004 | Yes |
| Q14B vs Qwen2.5-14B ZH | FPR | 17.1 | Yes | |
| Q14B vs Qwen2.5-14B EN | FPR | 48.0 | Yes | |
| QG vs GPT-4o-mini ZH | FPR | 4.2 | .041 | Yes |
| QG vs GPT-4o-mini EN | FPR | 24.0 | Yes | |
| McNemar’s test with continuity correction (conservative; without correction, | ||||
| all comparisons except Qwen3.5-9B vs. GPT EN IR remain significant at ); | ||||
| (IR), (FPR). | ||||
| QG: Qwen3.5-9BGPT-4o-mini cascade; Q14B: Qwen3.5-9BQwen2.5-14B cascade; GPT: GPT-4o-mini. | ||||
TC3 comparisons ( malicious, benign per language) are not included in Table 8; McNemar’s test on this subset remains future work due to the smaller sample size.
7.4.9. Comparison with Existing Defense Approaches
LGA differs from existing defense systems in two key dimensions. First, content-safety systems like Llama Guard (Inan et al., 2023) and NeMo Guardrails (Rebedea et al., 2023) operate at the text generation layer, filtering harmful model outputs before they reach the user. LGA operates one layer deeper, at the tool execution layer, intercepting unauthorized actions regardless of whether the generating text appears benign. A prompt injection that produces innocuous-looking text but triggers a malicious shell_exec call would evade content filters but be caught by LGA Layer 2.
Second, existing agent security benchmarks (InjecAgent (Zhan et al., 2024), AgentDojo (Debenedetti et al., 2024)) primarily measure attack success rates under various agent architectures, quantifying the problem without proposing a deployable mitigation. LGA provides both the architectural framework and empirical evidence for a specific defense mechanism (intent verification), bridging the gap between threat characterization and practical deployment. Moreover, the cascade architecture significantly reduces false positives below standalone cloud-judge performance (Table 8), offering a quantifiable usability advantage for deployment.
7.5. End-to-End Pipeline Evaluation
To validate that the four LGA layers operate in concert (RQ3), we run 100 samples (50 benign + 50 malicious, drawn using stratified sampling from the English TC1/TC2 subset) through the full pipeline: L1 (sandbox spawn via unshare), L2 (GPT-4o-mini intent judge), L3 (HMAC-SHA256 token signing + JSON schema validation), and L4 (append-only audit with fsync). Table 9 reports the results.
| Metric | Value |
| Interception Rate (IR) | 96.0% |
| False Positive Rate (FPR) | 16.0% |
| 89.6% | |
| Per-Layer Latency (ms) | |
| L1 (sandbox spawn) | P50 = 1.1, P99 = 3.4 |
| L2 (intent judge) | P50 = 966, P99 = 1622 |
| L3 (token + schema) | P50 0.1, P99 = 0.1c |
| L4 (audit + fsync) | P50 = 16.5, P99 = 17.6 |
| Total | P50 = 980, P99 = 1640 |
| Total latency percentiles are computed over the | |
| end-to-end distribution, not as the sum of per-layer percentiles. | |
| : harmonic mean of IR and ; see Sect. 7.1.4. | |
| cL3 P99 increase vs. microbenchmark (0.02 ms) reflects CPU scheduling jitter under concurrent pipeline execution. | |
The end-to-end pipeline achieves 96.0% IR with 16.0% FPR, consistent with standalone L2 performance (GPT-4o-mini: 95.4% IR, 12.3% FPR on the full English dataset). The FPR increase of +3.7 pp is within the expected sampling variance: with benign samples, the Clopper–Pearson 95% confidence interval for the observed 16% FPR spans , covering the full-dataset estimate of 12.3%. The wide CI reflects limited FPR precision at ; a larger-scale pipeline evaluation would be needed to precisely characterize end-to-end FPR. All blocks occur at Layer 2; Layers 1, 3, and 4 collectively add only 18 ms to total latency (P50), confirming that the non-judge layers impose negligible overhead. The dominant latency contributor remains the LLM judge (966 ms), reinforcing that LGA’s runtime cost is effectively the cost of a single inference call.
7.6. Generalization to External Benchmark
To assess whether our judges generalize beyond the synthetic benchmark (RQ1), we evaluate on InjecAgent (Zhan et al., 2024), an external indirect prompt injection benchmark with 1,054 test cases spanning 17 user tools and 62 attacker tools. All 1,054 InjecAgent cases are adversarial (no benign baseline); we uniformly sample 100 of these (fixed seed = 42; covering InjecAgent’s “direct harm” and “data stealing” categories—both involving indirect prompt injection via poisoned tool outputs, corresponding to TC1 indirect injection in our taxonomy), and convert them to our (task, tool_call) format by mapping each sample’s User Instruction field to the task description and the attacker’s injected tool_name/tool_parameters to the proposed tool call. Note that this conversion extracts the injected tool call in isolation, which may simplify detection compared to a full multi-turn pipeline where the injection is embedded in legitimate context; the resulting IR should therefore be interpreted as an upper bound under this evaluation protocol. We then run both GPT-4o-mini and Qwen3.5-9B judges with the same prompt template used in our main experiments.
| Judge | IR on InjecAgent | IR on Our Benchmark |
| (indirect subset, EN, ) | ||
| GPT-4o-mini | 100.0% | 95.6% |
| Qwen3.5-9B | 99.0% | 92.8% |
| EN indirect subset is the machine-translated | ||
| version of the ZH set (). | ||
Table 10 shows that both judges achieve near-perfect interception on InjecAgent (GPT-4o-mini: 100% [95% CI: 96.4–100%]; Qwen3.5-9B: 99% [95% CI: 94.6–100%]), matching or exceeding their performance on our synthetic indirect injection subset. This provides evidence that the intent verification approach generalizes to independently constructed attack scenarios and is not overfit to our benchmark’s generation process. Note that these results measure judge-level interception on isolated tool calls, not end-to-end agent compromise reduction; validating LGA’s impact on full-pipeline attack success rates remains future work.
8. Conclusion
The experimental findings in this paper support the architectural argument that governance boundaries—sandbox isolation, intent verification, zero-trust protocols, audit logging—are necessary complements to model capability improvements in autonomous agent deployments. We formalize three security threat classes in multi-agent systems—Agency Abuse, RAG Data Poisoning, and Malicious Skill Plugins—and propose LGA, a four-layer governance architecture with empirical evaluation across all three.
RQ1 (Intent verification effectiveness). Across four local LLM judges, two NLI baselines, and one cloud judge evaluated on a 1,081-sample bilingual benchmark, all five LLM-based judges achieve at least 93% overall interception on TC1/TC2, while lightweight NLI baselines remain below 10%. Qwen2.5-14B achieves the best local balance (98.2–98.5% IR, 9.7–20.1% FPR), approaching GPT-4o-mini (93.0–95.4% IR, 3.2–12.3% FPR). On TC3 (malicious skill plugins), interception among judges with meaningful precision–recall balance drops to 75–94%; judges at the high end of the raw range (e.g., Llama-3.1-8B ZH at 99.2%) achieve this via conservative blocking bias (87.5% FPR on benign plugin calls; see Sect. 7.4.3), suggesting that permission-boundary violations are harder to detect from intent alignment alone (; TC3 results warrant larger-scale validation), motivating complementary enforcement at Layers 1 and 3 (sandbox isolation and schema validation). Despite this TC3 limitation, generalization to the external InjecAgent benchmark yields 99–100% interception on indirect injection, providing preliminary evidence that the TC1/TC2 interception capability generalizes beyond our synthetic benchmark, subject to the evaluation protocol limitations noted above.
RQ2 (Security–latency trade-off). A two-stage cascade (Qwen3.5-9BGPT-4o-mini) achieves 91.9–92.6% IR with 1.9–6.7% FPR. For fully data-sovereign deployments, a local-only cascade (Qwen3.5-9BQwen2.5-14B) achieves 94.7–95.6% IR with 6.0–9.7% FPR (), indicating that a viable local alternative exists without cloud dependency. Qwen3.5-4B (3.4 GB, 490 ms) offers a viable edge-deployment option with 94.3–95.8% IR, though its elevated FPR (27.5–29.2%) limits suitability to risk-tolerant or human-in-the-loop settings, while Qwen2.5-14B approaches cloud-level precision on Chinese (FPR 9.7% vs. GPT-4o-mini’s 3.2%) at local-deployment cost.
RQ3 (End-to-end overhead). The full four-layer pipeline achieves 96% IR with a total P50 latency of 980 ms, of which Layers 1, 3, and 4 contribute only 18 ms. The LLM judge dominates runtime cost, confirming that LGA’s overhead is effectively one inference call.
Future work includes: (i) an open-source LGA reference implementation with optimized serving (vLLM/TensorRT) and risk-stratified tool-call routing; (ii) fine-tuning compact models on tool-call authorization datasets to close the NLI–LLM gap; (iii) evaluation under adaptive attacks, where adversaries iteratively optimize inputs against a known judge, to assess LGA’s robustness beyond the black-box threat model assumed here—our preliminary adversarial probe with purpose-built evasion samples (Sect. 7.4.5 (vii)) shows IR drops to 50–63% in a worst-case hand-crafted setting, motivating dedicated defenses; (iv) multilingual plugin permission verification, given the persistent cross-lingual TC3 asymmetry (Sect. 7.4.3) attributed to model-level Chinese reasoning differences—particularly for the Persistence subtype—and the open question of whether dedicated multilingual training can close this gap; (v) longitudinal study of how engineering responsibilities shift as AI capabilities mature.